OpenStudy (anonymous):

find all natural numbers x such that x^2+x+7 is a complete square

OpenStudy (anonymous):

x=0 as its undefined

OpenStudy (anonymous):

what?

OpenStudy (king):

@Diyadiya @Mimi_x3 @ParthKohli @lalaly check out this problem.....

OpenStudy (king):

1 is valid....

OpenStudy (anonymous):

make that 6

OpenStudy (king):

no hw 4...4^2+4+7=27...

OpenStudy (king):

yeah 6 is valid...

OpenStudy (king):

i think 1,6 are the only numbers but we have to prove that IF its true....

OpenStudy (king):

@satellite73 ....?

OpenStudy (anonymous):

hint: \[x^2<x^2+x+7<(x+3)^2\]

OpenStudy (king):

hw do we us e this? @Diyadiya @Mimi_x3 @apoorvk

OpenStudy (king):

@FoolForMath

OpenStudy (anonymous):

@King what are the complete squares between x^2 and (x+3)^2

OpenStudy (anonymous):

\[x^2<x^2+x+7<(x+3)^2 \\ then \ \ \ x^2+x+7=(x+1)^2 \ \ or \ \ x^2+x+7=(x+2)^2 \]

OpenStudy (king):

oh....

OpenStudy (king):

nice....:D